# Universal Solvability for Robot Motion Planning on Graphs

### v1

@@ 7.4 Augmenting a set of connected vertex and edges
Addition of vertices along with addition of edges to an existing graph $G(V,E)$ in order to make it universally solvable is also an interesting variant of the augmentation problem. Looking back at the instances showing lower bound of $\beta$ in 1-connected graphs, where a $p-1$ star required $\beta=\Theta(p)=\Theta(|V|)$ additional edges to make it universally solvable, we observe that adding two more leaves to the central vertex of the star graph makes it universally solvable (it is easy to verify that a star graph with $p+1$ leaves is universally solvable for $p$ robots). Therefore, instead of a linear number of edge additions ($\beta=\Theta(|V|)$), addition of vertices along with edges allow us to make the graph universally solvable by addition of just $2$ edges and $2$ vertices. This motivates us to look into a variation of the augmentation problem that takes into consideration the addition of both vertices and edges.
@@
We depict the graph $Z_{\alpha,\beta}$ in Figure 5. Therefore, the number of vertices $n$ of $Z_{\alpha,\beta}$ is equal to $1+(2\beta+1)\cdot(\alpha+1)$. We will show that $Z_{\alpha,\beta}$ for $n$ robots is a NO-instance of Graph Vertex and Edge Augmentation for Universal Solvability for vertex budget $\alpha$, and edge budget $\beta$.
@@ 8 Conclusion
Lastly, we consider the EAUS problem, where given a connected graph $G$ that is not universally solvable for $p$ robots, the question is to determine, if for a given $\beta$, at most $\beta$ edges can be added to $G$ to make it universally solvable for $p$ robots. Here, we analyse the problem based on the connectedness of the input graph. In particular, we provide an upper bound of $p-2$ as well as a lower bound of $\Theta(p)$ on $\beta$ for general graphs. Similarly, for the VEAUS problem, we find examples of graphs that require large enough vertices and edges to be added in order to make the graph universally solvable.
@@ A.1 Pseudocode of Algorithm 1
1:$(v_{1},\ldots,v_{n})\leftarrow$ BFS ordering from running BFS on $G$
2:$\mathfrak{T}\leftarrow$ BFS tree from running BFS on $G$
3:$V_{p}\leftarrow\{v_{1},\ldots,v_{p}\}$
@@
Notice that Algorithm 1 is a deterministic algorithm and the execution is only dependent on the set of vacant and occupied vertices (and not the exact robots occupying the vertices). Therefore, at each iteration of Algorithm 1, we have $S(i)=S^{\prime}(\pi(i))$. This is an invariant throughout the run of the Algorithm 1 and this completes the proof.
A.3 Pseudocode of Algorithm 2
@@
1:$n_{cc}\xleftarrow{}$ Number of connected components of $G$ (computed using DFS)
@@
7:return output of Algorithm $\mathcal{A}$ on input $(G(V,E),p,(S_{I},S_{R}))$triangleright Proposition 2.4
A.4 Proof of Theorem 5.3
@@
Given an instance $(G(V,E),p)$ for USolR, we state the deterministic algorithm, denoted as Algorithm 3, as follows. Algorithm 3 takes as input an instance $(G(V,E)\text{, }p)$, and first finds the connected components of the underlying graph $G$ using a Depth-First Search. If there are $2$ or more connected components in $G$, then Algorithm 3 returns that it is a NO-instance and terminates. Algorithm 3 then checks the reachability of $S_{\pi^{*}_{t}}$ from $S_{I}$ by invoking as a subroutine, Algorithm $\mathcal{A}$ (Proposition 2.4) for the FRMP instance $(G(V,E)\text{, }p\text{, }(S_{I},S_{\pi^{*}_{t}}))$ for all $t\in[p-1]$. If $\mathcal{A}$ outputs a NO for any of these $p-1$ instances (i.e., if there exists $t\in[p-1]\text{ such that }S_{\pi^{*}_{t}}$ is not reachable from $S_{I}$), then Algorithm 3 outputs that $G$ is not universally solvable for $p$ robots (i.e., $(G(V,E),p)$ is a NO-instance), else it outputs that $G$ is universally solvable for $p$ robots (i.e., $(G(V,E),p)$ is a YES-instance).
We formally describe Algorithm 3 below.
@@
1:$n_{cc}\xleftarrow{}$ Number of connected components of $G$ (computed using DFS)
@@
10:         $temp\leftarrow$ output of Algorithm $\mathcal{A}$ on input $(G(V,E),p,(S_{I},S_{\pi^{*}_{t}}))$triangleright Proposition 2.4
@@
The correctness of Algorithm 3 is immediate from Lemma 5.1. We now analyse the time complexity of Algorithm 3. Algorithm 3 takes as input a graph $G(V,E)$ and the number of robots $p$, and first performs a Depth-First Search to find the number of connected components of $G$, which takes at most linear time, $\mathcal{O}(|V|+|E|)$. Further, each of the $p-1$ reachability checks for $(S_{I},S_{\pi^{*}_{t}})$ for $t\in[p-1]$ by invoking algorithm $\mathcal{A}$, run in time $\mathcal{O}(|V|+|E|)$ (Proposition 2.4). Thus, the worst-case time complexity of Algorithm 3 is $\mathcal{O}((|V|+|E|)+(p-1)\cdot(|V|+|E|))=\mathcal{O}(p\cdot(|V|+|E|))$.
@@
Since the graph $G$ is connected, it can be decomposed into a tree of biconnected components (maximal biconnected subgraphs) called the block-cut tree [12] of the graph. The biconnected components are attached to each other at shared vertices that are cut vertices. Let $B_{g}=\{a_{1},a_{2},...,a_{b_{1}},a_{b_{1}+1},...,a_{b}\}$ be the set of biconnected components of $G$ such that $a_{1},a_{2},...,a_{b_{1}}$ are simple cycles, while $a_{b_{1}+1},a_{b_{1}+2}...,a_{b}$ are biconnected components that are not cycles.
@@
Next, we bound the sum of the sizes of the biconnected components. For this, we construct a rooted tree $T$ whose vertices are the biconnected components as follows: choose a biconnected component of $G$ that shares only one cut-vertex with other biconnected components. Let us denote this component as $a^{*}$. It is easy to prove that there exists at least one such biconnected component that shares only one cut-vertex with other biconnected components, as otherwise the entire graph would be just one biconnected component. Now, we consider the distance of this biconnected component to itself to be 0. For any other biconnected component $a^{\prime}$, we define its distance from $a^{*}$ to be the minimum number of biconnected components (including $a^{*}$) that need to be traversed to go from $a^{*}$ to $a^{\prime}$. For example, the biconnected components that share the single cut-vertex of $a^{*}$ are at a distance 1 from $a^{*}$ since we just need to traverse $a^{*}$ to reach them. Now, we construct the tree $T$ based on these distances, with the vertex corresponding to the component $a^{*}$ (with distance 00) as the root vertex. Then, the vertices corresponding to the distance-1 components are connected as children of the root. Similarly, the distance-2 components are connected as children of the distance-1 component that is adjacent to them, in the shortest path to $a^{*}$.
@@
Let $q$ be the number of cut vertices in the graph $G$. The number of edges in $T$ is $b-1$. Since each edge of the tree $T$ corresponds to exactly one cut vertex in $G$, the total number of times all cut vertices in $G$ are counted in the summation $\sum\limits_{i=1}^{d}s_{i}$ is $q+(b-1)$. Moreover, since vertices in $G$ that are not cut vertices are counted exactly once in the summation $\sum\limits_{i=1}^{d}s_{i}$, these are counted $(|V|-q)$ times in total. Adding these counts we get,
@@
Further, to obtain a bound on the number of biconnected components $b$, consider constructing the graph $G$ incrementally. We begin with a single vertex and iteratively add the remaining $|V|-1$ vertices along with their incident edges to incrementally create the biconnected components. Initially, with a single vertex, the count of biconnected components is zero. At each such step, adding a new vertex and connecting it to the existing graph can either expand an existing biconnected component or introduce at most one new biconnected component. Therefore, the total number of biconnected components is bounded above by the number of vertices minus one, i.e., $b\leq|V|-1$. Thus, we have,
@@
$<\frac{p}{2}\cdot\left(\sum_{i=1}^{b_{1}}s_{i}\right)+\frac{p}{2}\cdot\left(\sum_{i=b_{1}+1}^{b}s_{i}\right)$ (using Equation 1)
@@
$\leq\frac{p}{2}\cdot(2|V|-2)$ (using Equation 2)
@@ A.6 Pseudocode of Algorithm 4
1:$n_{cc}\leftarrow$ Number of connected components in $G$ (computed using DFS)
@@
4:else if $|E|\geq p\cdot|V|$ then $\triangleright$ Must be a YES-instance, by Section 6
@@
7:     return output of Algorithm 3 run on $G$ and $p$

### v2

@@ 7.4 Augmenting a set of connected vertex and edges
Addition of vertices along with addition of edges to an existing graph $G(V,E)$ in order to make it universally solvable is also an interesting variant of the augmentation problem. Looking back at the instances showing lower bound of $\beta$ in 1-connected graphs, where a p1p-1 star required $\beta=\Theta(p)=\Theta(|V|)$ additional edges to make it universally solvable, we observe that adding two more leaves to the central vertex of the star graph makes it universally solvable (it is easy to verify that a star graph with $p+1$ leaves is universally solvable for $p$ robots). Therefore, instead of a linear number of edge additions ($\beta=\Theta(|V|)$), addition of vertices along with edges allow us to make the graph universally solvable by addition of just $2$ edges and $2$ vertices. This motivates us to look into a variation of the augmentation problem that takes into consideration the addition of both vertices and edges.
@@
We depict the graph $Z_{\alpha,\beta}$ in Figure˜5. Therefore, the number of vertices $n$ of $Z_{\alpha,\beta}$ is equal to $1+(2\beta+1)\cdot(\alpha+1)$. We will show that $Z_{\alpha,\beta}$ for $n$ robots is a NO-instance of Graph Vertex and Edge Augmentation for Universal Solvability for vertex budget $\alpha$, and edge budget $\beta$.
@@ 8 Conclusion
Lastly, we consider the EAUS problem, where given a connected graph $G$ that is not universally solvable for $p$ robots, the question is to determine, if for a given $\beta$, at most $\beta$ edges can be added to $G$ to make it universally solvable for $p$ robots. Here, we analyse the problem based on the connectedness of the input graph. In particular, we provide an upper bound of p2p-2 as well as a lower bound of $\Theta(p)$ on $\beta$ for general graphs. Similarly, for the VEAUS problem, we find examples of graphs that require large enough vertices and edges to be added in order to make the graph universally solvable.
@@ A.1 Pseudocode of Algorithm˜1
[…]
@@
Notice that Algorithm˜1 is a deterministic algorithm and the execution is only dependent on the set of vacant and occupied vertices (and not the exact robots occupying the vertices). Therefore, at each iteration of Algorithm˜1, we have $S(i)=S^{\prime}(\pi(i))$. This is an invariant throughout the run of the Algorithm˜1 and this completes the proof.
A.3 Pseudocode of Algorithm˜2
@@
[…]
@@
7:return output of Algorithm $\mathcal{A}$ on input $(G(V,E),p,(S_{I},S_{R}))$ $\triangleright$ Proposition˜2.4
A.4 Proof of Theorem˜5.3
@@
Given an instance $(G(V,E),p)$ for USolR, we state the deterministic algorithm, denoted as Algorithm˜3, as follows. Algorithm˜3 takes as input an instance $(G(V,E)\text{, }p)$, and first finds the connected components of the underlying graph $G$ using a Depth-First Search. If there are $2$ or more connected components in $G$, then Algorithm˜3 returns that it is a NO-instance and terminates. Algorithm˜3 then checks the reachability of $S_{\pi^{*}_{t}}$ from $S_{I}$ by invoking as a subroutine, Algorithm $\mathcal{A}$ (Proposition˜2.4) for the FRMP instance $(G(V,E)\text{, }p\text{, }(S_{I},S_{\pi^{*}_{t}}))$ for all $t\in[p-1]$. If $\mathcal{A}$ outputs a NO for any of these p1p-1 instances (i.e., if there exists $t\in[p-1]\text{ such that }S_{\pi^{*}_{t}}$ is not reachable from $S_{I}$), then Algorithm˜3 outputs that $G$ is not universally solvable for $p$ robots (i.e., $(G(V,E),p)$ is a NO-instance), else it outputs that $G$ is universally solvable for $p$ robots (i.e., $(G(V,E),p)$ is a YES-instance).
We formally describe Algorithm˜3 below.
@@
[…]
@@
10:   $temp\leftarrow$ output of Algorithm $\mathcal{A}$ on input $(G(V,E),p,(S_{I},S_{\pi^{*}_{t}}))$ $\triangleright$ Proposition˜2.4
@@
The correctness of Algorithm˜3 is immediate from Lemma˜5.1. We now analyse the time complexity of Algorithm˜3. Algorithm˜3 takes as input a graph $G(V,E)$ and the number of robots $p$, and first performs a Depth-First Search to find the number of connected components of $G$, which takes at most linear time, $\mathcal{O}(|V|+|E|)$. Further, each of the p1p-1 reachability checks for $(S_{I},S_{\pi^{*}_{t}})$ for $t\in[p-1]$ by invoking algorithm $\mathcal{A}$, run in time $\mathcal{O}(|V|+|E|)$ (Proposition˜2.4). Thus, the worst-case time complexity of Algorithm˜3 is $\mathcal{O}((|V|+|E|)+(p-1)\cdot(|V|+|E|))=\mathcal{O}(p\cdot(|V|+|E|))$.
@@
Since the graph $G$ is connected, it can be decomposed into a tree of biconnected components (maximal biconnected subgraphs) called the block-cut tree [10.1145/362248.362272] of the graph. The biconnected components are attached to each other at shared vertices that are cut vertices. Let $B_{g}=\{a_{1},a_{2},...,a_{b_{1}},a_{b_{1}+1},...,a_{b}\}$ be the set of biconnected components of $G$ such that $a_{1},a_{2},...,a_{b_{1}}$ are simple cycles, while $a_{b_{1}+1},a_{b_{1}+2}...,a_{b}$ are biconnected components that are not cycles.
@@
Next, we bound the sum of the sizes of the biconnected components. For this, we construct a rooted tree $T$ whose vertices are the biconnected components as follows: choose a biconnected component of $G$ that shares only one cut-vertex with other biconnected components. Let us denote this component as $a^{*}$. It is easy to prove that there exists at least one such biconnected component that shares only one cut-vertex with other biconnected components, as otherwise the entire graph would be just one biconnected component. Now, we consider the distance of this biconnected component to itself to be 0. For any other biconnected component $a^{\prime}$, we define its distance from $a^{*}$ to be the minimum number of biconnected components (including $a^{*}$) that need to be traversed to go from $a^{*}$ to $a^{\prime}$. For example, the biconnected components that share the single cut-vertex of $a^{*}$ are at a distance 1 from $a^{*}$ since we just need to traverse $a^{*}$ to reach them. Now, we construct the tree $T$ based on these distances, with the vertex corresponding to the component $a^{*}$ (with distance 0) as the root vertex. Then, the vertices corresponding to the distance-1 components are connected as children of the root. Similarly, the distance-2 components are connected as children of the distance-1 component that is adjacent to them, in the shortest path to $a^{*}$.
@@
Let $q$ be the number of cut vertices in the graph $G$. The number of edges in $T$ is b1b-1. Since each edge of the tree $T$ corresponds to exactly one cut vertex in $G$, the total number of times all cut vertices in $G$ are counted in the summation $\sum\limits_{i=1}^{d}s_{i}$ is q+(b1)q+(b-1). Moreover, since vertices in $G$ that are not cut vertices are counted exactly once in the summation $\sum\limits_{i=1}^{d}s_{i}$, these are counted (|V|q)(|V|-q) times in total. Adding these counts we get,
@@
Further, to obtain a bound on the number of biconnected components $b$, consider constructing the graph $G$ incrementally. We begin with a single vertex and iteratively add the remaining |V|1|V|-1 vertices along with their incident edges to incrementally create the biconnected components. Initially, with a single vertex, the count of biconnected components is zero. At each such step, adding a new vertex and connecting it to the existing graph can either expand an existing biconnected component or introduce at most one new biconnected component. Therefore, the total number of biconnected components is bounded above by the number of vertices minus one, i.e., $b\leq|V|-1$. Thus, we have,
@@
$<\frac{p}{2}\cdot\left(\sum_{i=1}^{b_{1}}s_{i}\right)+\frac{p}{2}\cdot\left(\sum_{i=b_{1}+1}^{b}s_{i}\right)$ (using Equation˜1)
@@
$\leq\frac{p}{2}\cdot(2|V|-2)$ (using Equation˜2)
@@ A.6 Pseudocode of Algorithm˜4
[…]
@@
4:else if $|E|\geq p\cdot|V|$ then $\triangleright$ Must be a YES-instance, by Section˜6
@@
7:  return output of Algorithm˜3 run on $G$ and $p$